OpenStudy (anonymous):

an=12an-1. the first term is 0.001. which is the explicit formula to find the nth term. a. 0.001*12^n b.12*0.001^n c.12*0.001^n-1 d.0.001*12^n+1 e.0.001*12^n-1

OpenStudy (anonymous):

Let a=.001 \[a_2= 12 a\\ a_3= 12 a_2 = 12^2 a\\ a_4= 12 a_3 = 12^3 a\\ \]

OpenStudy (anonymous):

\[a_n=12^{n-1} a= .001 12^{n-1}\]